# One-Prompt Strikes Back: Sparse Mixture of Experts for Prompt-based Continual Learning

### v1

@@ Abstract
Prompt-based methods have recently gained prominence in Continual Learning (CL) due to their strong performance and memory efficiency. A prevalent strategy in this paradigm assigns a dedicated subset of prompts to each task, which, while effective, incurs substantial computational overhead and causes memory requirements to scale linearly with the number of tasks. Conversely, approaches employing a single shared prompt across tasks offer greater efficiency but often suffer from degraded performance due to knowledge interference. To reconcile this trade-off, we propose SMoPE, a novel framework that integrates the benefits of both task-specific and shared prompt strategies. Inspired by recent findings on the relationship between Prefix Tuning and Mixture of Experts (MoE), SMoPE organizes a shared prompt into multiple "prompt experts" within a sparse MoE architecture. For each input, only a select subset of relevant experts is activated, effectively mitigating interference. To facilitate expert selection, we introduce a prompt-attention score aggregation mechanism that computes a unified proxy score for each expert, enabling dynamic and sparse activation. Additionally, we propose an adaptive noise mechanism to encourage balanced expert utilization while preserving knowledge from prior tasks. To further enhance expert specialization, we design a prototype-based loss function that leverages prefix keys as implicit memory representations. Extensive experiments across multiple CL benchmarks demonstrate that SMoPE consistently outperforms task-specific prompt methods and achieves performance competitive with state-of-the-art approaches, all while significantly reducing parameter counts and computational costs.
@@ 1 Introduction
Continual Learning (CL) is a critical research area focused on enabling neural networks to learn from a sequence of tasks in dynamic environments while retaining knowledge from previous tasks (aljundi2017expert, ; de2021continual, ; zhang2023slca, ). A primary challenge in CL is catastrophic forgetting, where performance on earlier tasks deteriorates as new ones are learned (mccloskey1989catastrophic, ; french1999catastrophic, ; mehta2023empirical, ). Recently, prompt-based approaches have gained attention as a promising direction in CL, offering strong performance and high memory efficiency (wang2022learning, ; wang2022dualprompt, ; smith2023coda, ; le2024mixture, ). These methods adapt pre-trained models using a small set of learnable parameters, referred to as prompts, which function as task-specific instructions to guide adaptation and alleviate forgetting. Several studies have demonstrated the effectiveness of prompt-based methods, reporting state-of-the-art results across a range of CL benchmarks.
A common strategy in prompt-based CL is to allocate a distinct subset of prompt parameters to each task (wang2022dualprompt, ; wang2023hierarchical, ; le2024mixture, ). This task-specific partitioning helps mitigate interference by isolating knowledge within separate prompt modules. While effective, such approaches face several notable limitations. First, when task identity is unknown at inference time, the model must infer the appropriate prompt for each input. Existing methods often rely on forwarding the input through the full pre-trained model to compute a query, introducing non-negligible computational overhead (huang2024ovor, ; kim2024one, ). Second, assigning dedicated prompts to each task hinders scalability and limits knowledge sharing. As the number of tasks increases, the number of learnable prompt parameters grows linearly, making this approach inefficient for long-term continual learning. Moreover, strict prompt partitioning restricts transferability by preventing the reuse or adaptation of previously learned prompts, thereby limiting positive knowledge transfer across tasks.
In contrast to task-specific prompt methods, huang2024ovor recently introduced OVOR, a highly parameter-efficient and computationally lightweight approach that employs a single shared prompt across all tasks. Despite its efficiency, OVOR underperforms relative to state-of-the-art task-specific prompt methods. We hypothesize that this performance gap arises from excessive knowledge interference: because the same prompt is continually updated across tasks, it struggles to retain task-specific information, leading to degraded performance. This raises a key question:
@@
Most prompt-based CL methods rely on Prefix Tuning (li2021prefix, ) to integrate prompts into pre-trained models. However, recent work by le2024mixture ; le2025expressiveness offers a new perspective: each attention head can be viewed as a composition of multiple Mixture of Experts (MoE) models (jacobs1991adaptive, ; shazeer2017outrageously, ), and prefix tuning effectively introduces new prompt experts into these models. Building on this insight, we propose SMoPE (Sparse Mixture of Prompt Experts), a novel method that retains a single shared prompt while structuring it as multiple prompt experts within a sparse MoE framework.
@@ 2.1 Continual Learning
Continual Learning (CL) involves training a model on a sequence of tasks $\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$. Each task $\mathcal{D}_{t}=\{({\bm{x}}_{i}^{t},y_{i}^{t})\}_{i=1}^{\mathcal{N}_{t}}$ contains $\mathcal{N}_{t}$ samples, where ${\bm{x}}_{i}^{t}\in\mathcal{X}^{t}$ is an input and $y_{i}^{t}\in\mathcal{Y}^{t}$ is its corresponding label. The primary objective is for the model, after being trained on task $t$, to perform well on the set of all classes encountered up to task $t$, denoted by $\mathcal{Y}^{1:t}=\bigcup_{i=1}^{t}\mathcal{Y}^{i}$. We focus on the challenging class-incremental learning scenario (van2019three, ; wang2022learning, ; wang2023hierarchical, ), where the label spaces of different tasks are disjoint, i.e., $\mathcal{Y}^{t}\bigcap\mathcal{Y}^{t^{\prime}}=\varnothing$ for any $t\neq t^{\prime}$. A standard constraint in CL is that data from previous tasks $\mathcal{D}_{1},\dots,\mathcal{D}_{t-1}$ is unavailable when training on the current task $\mathcal{D}_{t}$. This sequential training process, without access to past data, makes the model susceptible to catastrophic forgetting (mccloskey1989catastrophic, ; nguyen2019toward, ; mehta2023empirical, ), wherein performance on previously learned tasks degrades significantly as the model adapts to new ones.
@@ 2.2 Prompt-based Continual Learning
In vision tasks, prompting is typically applied to the Vision Transformer (ViT) (dosovitskiy2020image, ), which consists of a sequence of Multi-Head Self-Attention (MSA) (vaswani2017attention, ) blocks. To illustrate how prompts are integrated, we first describe the standard MSA mechanism. Let the input to an MSA layer be a sequence of token embeddings ${\bm{x}}_{1},\dots,{\bm{x}}_{N}]^{\top}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length and $d$ is the embedding dimension. The MSA operation is defined as follows:
@@
Most prompt-based methods (wang2022dualprompt, ; wang2023hierarchical, ; jiao2024vector, ) adopt Prefix Tuning (li2021prefix, ), which introduces learnable prefix key ${\bm{P}}^{K}\in\mathbb{R}^{N_{p}\times d}$ and prefix value ${\bm{P}}^{V}\in\mathbb{R}^{N_{p}\times d}$ parameters, prepended to the input key and value matrices of the MSA layer:
@@
Conventional prompt-based continual learning methods, such as DualPrompt (wang2022dualprompt, ), HiDe-Prompt (wang2023hierarchical, ), and NoRGa (le2024mixture, ), typically allocate a separate set of prompt parameters for each task. However, recent work by huang2024ovor demonstrates that a single prompt shared across tasks can achieve performance comparable to that of task-specific prompting, raising questions about the efficiency and utilization of prompts in prior approaches.
@@ 2.3 Mixture of Experts
Mixture of Experts (MoE) extends classical mixture models by introducing an adaptive gating mechanism (jacobs1991adaptive, ; jordan1994hierarchical, ). For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model computes outputs from $N^{\prime}$ experts $f_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$, which are then combined using weights from a learned gating function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ as follows:
@@
where $s_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denotes the score function for expert $f_{j}$. To enhance scalability, shazeer2017outrageously introduced the Sparse Mixture of Experts (SMoE), a variant that activates only the top-$K$ experts with the highest scores. The set of these indices, denoted $K_{\bm{h}}$, is formally defined as:
@@
By routing only a small subset of experts per input, SMoE achieves high model capacity with limited computational overhead. This property has driven adoption in large-scale systems, including language models (fedus2022switch, ; jiang2024mixtral, ; comanici2025gemini, ), computer vision (riquelme2021scaling, ; xue2023raphael, ), and multi-task learning (ma2018modeling, ; yang2024multi, ).
@@ 3.1 Mixture of Experts and Prefix Tuning
We adopt prefix tuning as our primary prompting strategy, following prior work. Furthermore, we use a single prompt shared across all tasks, similar to huang2024ovor . Recent findings by le2024mixture show that each attention head in a ViT can be viewed as a structured composition of multiple MoE models, framing prefix tuning as a way to add new experts to this structure.
@@ 3.2 Sparse Mixture of Prompt Experts
Similar to OVOR (huang2024ovor, ), SMoPE employs a single prompt shared across all tasks, ensuring high parameter efficiency. However, unlike OVOR, SMoPE activates only a sparse subset of $K$ prompt experts per input $\bm{X}$} within each MSA layer, introducing implicit parameter partitioning. By updating only the relevant parameters, SMoPE reduces knowledge interference and mitigates catastrophic forgetting. Importantly, this parameter selection relies solely on the current layer’s input $\bm{X}$} and does not require precomputed query representations from a full model forward pass, as in prior task-specific prompting methods (wang2022dualprompt, ; wang2023hierarchical, ; smith2023coda, ; le2024mixture, ). As a result, SMoPE offers substantial reductions in computational overhead compared to these approaches (kim2024one, ). Furthermore, by dynamically reusing relevant experts across tasks, SMoPE naturally promotes knowledge sharing and transfer.
@@ 3.3 Balancing Expert Utilization with Adaptive Noise
A well-known challenge in training SMoE models is the imbalance in expert utilization, where a small subset of experts dominates routing decisions while others remain underutilized or inactive (mu2025comprehensive, ). This imbalance reduces the diversity of learned representations and ultimately limits model effectiveness. We find that SMoPE exhibits a similar issue: during training, only a small group of prompt experts is consistently activated across tasks. While SMoPE mitigates interference by selectively activating experts, repeatedly relying on the same subset still risks knowledge interference.
@@ 3.4 Prefix Key Prototypes for Expert Specialization
Additional Training Strategies. In continual learning, the MLP classifier head often develops a bias toward newly introduced classes (hou2019learning, ; belouadah2019il2m, ). To mitigate this, we adopt task-adaptive prediction, following prior work (wang2023hierarchical, ; jiao2024vector, ; le2024mixture, ). This technique adjusts predictions using the representation statistics of previously learned classes to correct classifier bias and stabilize prompt learning. Additionally, inspired by advances in SMoE training (wu2022residual, ; lin2024moe, ; cai2025survey, ), we apply dense expert training (i.e., without sparse selection) during the initial epochs of the first task. This aids in establishing stable expert representations before enabling sparse routing.
@@ 4.1 Experimental Details
Datasets. Following jiao2024vector , we evaluate SMoPE on three representative CL benchmarks: ImageNet-R (boschini2022transfer, ), CIFAR-100 (krizhevsky2009learning, ), and CUB-200 (wah2011caltech, ). ImageNet-R consists of 200 challenging classes, including difficult samples from ImageNet and newly collected data with diverse stylistic variations, and is split into 5, 10, or 20 disjoint tasks. CIFAR-100 contains 100 classes, which we randomly partition into 10 tasks. CUB-200 comprises fine-grained images of 200 bird species, randomly divided into 10 tasks with 20 classes each.
@@
Baselines. We compare our approach against several representative prompt-based methods, including L2P (wang2022learning, ), DualPrompt (wang2022dualprompt, ), CODA-Prompt (smith2023coda, ), and VQ-Prompt (jiao2024vector, ). We also consider OVOR (huang2024ovor, ), which employs a single prompt, and state-of-the-art task-specific prompt methods such as HiDe-Prompt (wang2023hierarchical, ) and NoRGa (le2024mixture, ). For L2P, we follow smith2023coda and include two variants: "L2P++", which replaces prompt tuning with prefix tuning, and "Deep L2P++", which extends L2P++ by inserting prompts into the first five MSA blocks. Finally, we report the performance of "Joint-Train", an offline upper bound obtained by training on all tasks simultaneously.
Implementation Details. All experiments are conducted on a single NVIDIA A100 GPU. To ensure a fair comparison, all baseline methods are re-implemented using the configurations reported in their original papers. Following jiao2024vector , we adopt a ViT-B/16 backbone pretrained on ImageNet-1K (russakovsky2015imagenet, ) and ImageNet-21K (ridnik2021imagenet, ) as the backbone. We also evaluate self-supervised ViT-B/16 models from iBOT-1K (zhou2021ibot, ) and DINO-1K (caron2021emerging, ). The prompt length is fixed at $N_{p}=25$, and the number of selected prompt experts is set to $K=5$ for all experiments. Prompts are inserted into the first six MSA blocks. SMoPE is optimized using AdamW (loshchilov2017decoupled, ) with a cosine learning rate decay schedule. The batch size is set to 64 for ImageNet-R, and 128 for both CIFAR-100 and CUB-200. Additional implementation details are provided in Appendix C.

### v2

@@ Abstract
Prompt-based methods have recently gained prominence in Continual Learning (CL) due to their strong performance and memory efficiency. A prevalent strategy in this paradigm assigns a dedicated subset of prompts to each task, which, while effective, incurs substantial computational overhead and causes memory requirements to scale linearly with the number of tasks. Conversely, approaches employing a single shared prompt across tasks offer greater efficiency but often suffer from degraded performance due to knowledge interference. To reconcile this trade-off, we propose SMoPE, a novel framework that integrates the benefits of both task-specific and shared prompt strategies. Inspired by recent findings on the relationship between Prefix Tuning and Mixture of Experts (MoE), SMoPE organizes a shared prompt into multiple "prompt experts" within a sparse MoE architecture. For each input, only a select subset of relevant experts is activated, effectively mitigating interference. To facilitate expert selection, we introduce a prompt-attention score aggregation mechanism that computes a unified proxy score for each expert, enabling dynamic and sparse activation. Additionally, we propose an adaptive noise mechanism to encourage balanced expert utilization while preserving knowledge from prior tasks. To further enhance expert specialization, we design a prototype-based loss function that leverages prefix keys as implicit memory representations. Extensive experiments across multiple CL benchmarks demonstrate that SMoPE consistently outperforms task-specific prompt methods and achieves performance competitive with state-of-the-art approaches, all while significantly reducing parameter counts and computational costs. Our code is publicly available at https://github.com/Minhchuyentoancbn/SMoPE.
@@ 1 Introduction
Continual Learning (CL) is a critical research area focused on enabling neural networks to learn from a sequence of tasks in dynamic environments while retaining knowledge from previous tasks (Aljundi et al., 2017; De Lange et al., 2021; Zhang et al., 2023). A primary challenge in CL is catastrophic forgetting, where performance on earlier tasks deteriorates as new ones are learned (McCloskey & Cohen, 1989; French, 1999; Mehta et al., 2023). Recently, prompt-based approaches have gained attention as a promising direction in CL, offering strong performance and high memory efficiency (Wang et al., 2022b; a; Smith et al., 2023; Le et al., 2024). These methods adapt pre-trained models using a small set of learnable parameters, referred to as prompts, which function as task-specific instructions to guide adaptation and alleviate forgetting. Several studies have demonstrated the effectiveness of prompt-based methods, reporting state-of-the-art results across a range of CL benchmarks.
A common strategy in prompt-based CL is to allocate a distinct subset of prompt parameters to each task (Wang et al., 2022a; 2023; Le et al., 2024). This task-specific partitioning helps mitigate interference by isolating knowledge within separate prompt modules. While effective, such approaches face several notable limitations. First, when task identity is unknown at inference time, the model must infer the appropriate prompt for each input. Existing methods often rely on forwarding the input through the full pre-trained model to compute a query, introducing non-negligible computational overhead (Huang et al., 2024; Kim et al., 2024). Second, assigning dedicated prompts to each task hinders scalability and limits knowledge sharing. As the number of tasks increases, the number of learnable prompt parameters grows linearly, making this approach inefficient for long-term continual learning. Moreover, strict prompt partitioning restricts transferability by preventing the reuse or adaptation of previously learned prompts, thereby limiting positive knowledge transfer across tasks.
In contrast to task-specific prompt methods, Huang et al. (2024) recently introduced OVOR, a highly parameter-efficient and computationally lightweight approach that employs a single shared prompt across all tasks. Despite its efficiency, OVOR underperforms relative to state-of-the-art task-specific prompt methods. We hypothesize that this performance gap arises from excessive knowledge interference: because the same prompt is continually updated across tasks, it struggles to retain task-specific information, leading to degraded performance. This raises a key question:
@@
Most prompt-based CL methods rely on Prefix Tuning (Li & Liang, 2021) to integrate prompts into pre-trained models. However, recent work by Le et al. (2024) offers a new perspective: each attention head can be viewed as a composition of multiple Mixture of Experts (MoE) models (Jacobs et al., 1991; Shazeer et al., 2017), and prefix tuning effectively introduces new prompt experts into these models. Building on this insight, we propose SMoPE (Sparse Mixture of Prompt Experts), a novel method that retains a single shared prompt while structuring it as multiple prompt experts within a sparse MoE framework.
@@ 2.1 Continual Learning
Continual Learning (CL) involves training a model on a sequence of tasks $\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$. Each task $\mathcal{D}_{t}=\{({\bm{x}}_{i}^{t},y_{i}^{t})\}_{i=1}^{\mathcal{N}_{t}}$ contains $\mathcal{N}_{t}$ samples, where ${\bm{x}}_{i}^{t}\in\mathcal{X}^{t}$ is an input and $y_{i}^{t}\in\mathcal{Y}^{t}$ is its corresponding label. The primary objective is for the model, after being trained on task $t$, to perform well on the set of all classes encountered up to task $t$, denoted by $\mathcal{Y}^{1:t}=\bigcup_{i=1}^{t}\mathcal{Y}^{i}$. We focus on the challenging class-incremental learning scenario (Van de Ven & Tolias, 2019; Wang et al., 2022b; 2023), where the label spaces of different tasks are disjoint, i.e., $\mathcal{Y}^{t}\bigcap\mathcal{Y}^{t^{\prime}}=\varnothing$ for any $t\neq t^{\prime}$. A standard constraint in CL is that data from previous tasks $\mathcal{D}_{1},\dots,\mathcal{D}_{t-1}$ is unavailable when training on the current task $\mathcal{D}_{t}$. This sequential training process, without access to past data, makes the model susceptible to catastrophic forgetting (McCloskey & Cohen, 1989; Nguyen et al., 2019; Mehta et al., 2023), wherein performance on previously learned tasks degrades significantly as the model adapts to new ones.
@@ 2.2 Prompt-based Continual Learning
In vision tasks, prompting is typically applied to the Vision Transformer (ViT) (Dosovitskiy et al., 2020), which consists of a sequence of Multi-Head Self-Attention (MSA) (Vaswani et al., 2017) blocks. To illustrate how prompts are integrated, we first describe the standard MSA mechanism. Let the input to an MSA layer be a sequence of token embeddings ${\bm{x}}_{1},\dots,{\bm{x}}_{N}]^{\top}\in\mathbb{R}^{N\times d}$, where $N$ is the sequence length and $d$ is the embedding dimension. The MSA operation is defined as follows:
@@
Most prompt-based methods (Wang et al., 2022a; 2023; Jiao et al., 2024) adopt Prefix Tuning (Li & Liang, 2021), which introduces learnable prefix key ${\bm{P}}^{K}\in\mathbb{R}^{N_{p}\times d}$ and prefix value ${\bm{P}}^{V}\in\mathbb{R}^{N_{p}\times d}$ parameters, prepended to the input key and value matrices of the MSA layer:
@@
Conventional prompt-based continual learning methods, such as DualPrompt (Wang et al., 2022a), HiDe-Prompt (Wang et al., 2023), and NoRGa (Le et al., 2024), typically allocate a separate set of prompt parameters for each task. However, recent work by Huang et al. (2024) demonstrates that a single prompt shared across tasks can achieve performance comparable to that of task-specific prompting, raising questions about the efficiency and utilization of prompts in prior approaches. While subsequent works have attempted to address related issues, they still present limitations. For instance, the mechanism proposed by Gao et al. (2024) to mitigate incorrect prompt selection at test time still results in a linear growth of prompts with the number of tasks. Similarly, the strategy of Roy et al. (2024) to manage prompt growth by employing Large Language Models (LLMs) introduces significant computational overhead due to its reliance on an external model.
@@ 2.3 Mixture of Experts
Mixture of Experts (MoE) extends classical mixture models by introducing an adaptive gating mechanism (Jacobs et al., 1991; Jordan & Jacobs, 1994). For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model computes outputs from $N^{\prime}$ experts $f_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$, which are then combined using weights from a learned gating function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ as follows:
@@
where $s_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denotes the score function for expert $f_{j}$. To enhance scalability, Shazeer et al. (2017) introduced the Sparse Mixture of Experts (SMoE), a variant that activates only the top-$K$ experts with the highest scores. The set of these indices, denoted $K_{\bm{h}}$, is formally defined as:
@@
By routing only a small subset of experts per input, SMoE achieves high model capacity with limited computational overhead. This property has driven adoption in large-scale systems, including language models (Fedus et al., 2022; Jiang et al., 2024; Comanici et al., 2025), computer vision (Riquelme et al., 2021; Xue et al., 2023), and multi-task learning (Ma et al., 2018; Yang et al., 2024).
@@ 3.1 Mixture of Experts and Prefix Tuning
We adopt prefix tuning as our primary prompting strategy, following prior work. Furthermore, we use a single prompt shared across all tasks, similar to Huang et al. (2024). Recent findings by Le et al. (2024) show that each attention head in a ViT can be viewed as a structured composition of multiple MoE models, framing prefix tuning as a way to add new experts to this structure.
@@ 3.2 Sparse Mixture of Prompt Experts
Similar to OVOR (Huang et al., 2024), SMoPE employs a single prompt shared across all tasks, ensuring high parameter efficiency. However, unlike OVOR, SMoPE activates only a sparse subset of $K$ prompt experts per input $\bm{X}$} within each MSA layer, introducing implicit parameter partitioning. By updating only the relevant parameters, SMoPE reduces knowledge interference and mitigates catastrophic forgetting. Importantly, this parameter selection relies solely on the current layer’s input $\bm{X}$} and does not require precomputed query representations from a full model forward pass, as in prior task-specific prompting methods (Wang et al., 2022a; 2023; Smith et al., 2023; Le et al., 2024). As a result, SMoPE offers substantial reductions in computational overhead compared to these approaches (Kim et al., 2024). Furthermore, by dynamically reusing relevant experts across tasks, SMoPE naturally promotes knowledge sharing and transfer.
@@ 3.3 Balancing Expert Utilization with Adaptive Noise
A well-known challenge in training SMoE models is the imbalance in expert utilization, where a small subset of experts dominates routing decisions while others remain underutilized or inactive (Mu & Lin, 2025). This imbalance reduces the diversity of learned representations and ultimately limits model effectiveness. We find that SMoPE exhibits a similar issue: during training, only a small group of prompt experts is consistently activated across tasks. While SMoPE mitigates interference by selectively activating experts, repeatedly relying on the same subset still risks knowledge interference.
@@ 3.4 Prefix Key Prototypes for Expert Specialization
Additional Training Strategies. In continual learning, the MLP classifier head often develops a bias toward newly introduced classes (Hou et al., 2019; Belouadah & Popescu, 2019). To mitigate this, we adopt task-adaptive prediction, following prior work (Wang et al., 2023; Jiao et al., 2024; Le et al., 2024). This technique adjusts predictions using the representation statistics of previously learned classes to correct classifier bias and stabilize prompt learning. Additionally, inspired by advances in SMoE training (Wu et al., 2022; Lin et al., 2024; Cai et al., 2025), we apply dense expert training (i.e., without sparse selection) during the initial epochs of the first task. This aids in establishing stable expert representations before enabling sparse routing.
@@ 4.1 Experimental Details
Datasets. Following Jiao et al. (2024), we evaluate SMoPE on three representative CL benchmarks: ImageNet-R (Boschini et al., 2022), CIFAR-100 (Krizhevsky et al., 2009), and CUB-200 (Wah et al., 2011). ImageNet-R consists of 200 challenging classes, including difficult samples from ImageNet and newly collected data with diverse stylistic variations, and is split into 5, 10, or 20 disjoint tasks. CIFAR-100 contains 100 classes, which we randomly partition into 10 tasks. CUB-200 comprises fine-grained images of 200 bird species, randomly divided into 10 tasks with 20 classes each.
@@
Baselines. We compare our approach against several representative prompt-based continual learning methods, including L2P (Wang et al., 2022b), DualPrompt (Wang et al., 2022a) and CODA-Prompt (Smith et al., 2023). Additionally, we evaluate against recent state-of-the-art methods, such as ConvPrompt (Roy et al., 2024), CPrompt (Gao et al., 2024), and VQ-Prompt (Jiao et al., 2024). We also consider OVOR (Huang et al., 2024), which employs a single prompt, and state-of-the-art task-specific prompt methods such as HiDe-Prompt (Wang et al., 2023) and NoRGa (Le et al., 2024). For L2P, we follow Smith et al. (2023) and include two variants: "L2P++", which replaces prompt tuning with prefix tuning, and "Deep L2P++", which extends L2P++ by inserting prompts into the first five MSA blocks. Finally, we report the performance of "Joint-Train", an offline upper bound obtained by training on all tasks simultaneously.
Implementation Details. All experiments are conducted on a single NVIDIA A100 GPU. To ensure a fair comparison, all baseline methods are re-implemented using the configurations reported in their original papers. Following Jiao et al. (2024), we adopt a ViT-B/16 backbone pretrained on ImageNet-1K (Russakovsky et al., 2015) and ImageNet-21K (Ridnik et al., 2021) as the backbone. We also evaluate self-supervised ViT-B/16 models from iBOT-1K (Zhou et al., 2021) and DINO-1K (Caron et al., 2021). The prompt length is fixed at $N_{p}=25$, and the number of selected prompt experts is set to $K=5$ for all experiments. Prompts are inserted into the first six MSA blocks. SMoPE is optimized using AdamW (Loshchilov & Hutter, 2017) with a cosine learning rate decay schedule. The batch size is set to 64 for ImageNet-R, and 128 for both CIFAR-100 and CUB-200. Additional implementation details are provided in Appendix C.